### v1

# Negative feedback and oscillations in a model for mRNA translation

Abstract
The ribosome flow model (RFM) is a phenomenological model for the unidirectional flow of particles along a 1D chain of $n$ sites. The RFM has been extensively used to study the dynamics of ribosome flow along a single mRNA molecule during translation. In this case, the particles model ribosomes and each site corresponds to a consecutive group of codons. Networks of interconnected RFMs have been used to model and analyze large-scale translation in the cell and, in particular, the effects of competition for shared resources. Here, we analyze the RFM with a negative feedback connection from the protein production rate to the initiation rate. This models, for example, the production of proteins that inhibit the translation of their own mRNA. Using tools from the theory of 2-cooperative dynamical systems, we provide a simple condition guaranteeing that the closed-loop system admits at least one non-trivial periodic solution. When this condition holds, we also explicitly characterize a large set of initial conditions such that any solution emanating from this set converges to a non-trivial periodic solution. Such a solution corresponds to a periodic pattern of ribosome densities along the mRNA, and to a periodic pattern of protein production.
Index Terms:
mRNA translation, periodic solutions, regulation of gene expression.
I Introduction
Living cells produce proteins using RNA molecules as templates. Three consecutive nucleotides in the mRNA form a codon that encodes for a specific amino-acid. Complex molecular machines, called ribosomes, scan the mRNA sequentially, codon by codon, and generate a chain of corresponding amino-acids that eventually becomes the coded protein. Every ribosome that reaches the stop codon releases the chain of amino acids, so the rate of ribosomes leaving the chain is the protein production rate. To pipeline this process, several ribosomes may decode the same mRNA molecule simultaneously, thus generating a polysome.
The dynamics of ribosome flow along the mRNA is determined by various bio-physical properties including the mRNA sequence, abundance of decoding factors, and the interactions between ribosomes. Indeed, since every ribosome must follow the mRNA in a sequential manner, a slow-moving ribosome may lead to the generation of “traffic jams” of ribosomes behind it. Such traffic jams have been implicated with various diseases including Huntington disease [40],  [39].
A fundamental problem in biology and medicine is understanding the regulation and dynamics of ribosome flow and the protein production rate. Rigorous analysis of ribosome flow is also important in scientific fields that include natural or artificial manipulations of the translation machinery such as biotechnology, synthetic biology, and mRNA viruses.
Computational modeling of ribosome flow is attracting considerable interest [44, 7, 51]. This is motivated also by new experimental methods, like ribosome profiling, that provide considerable information on which mRNAs are being actively translated in a cell at a given time. The totally asymmetric simple exclusion process (TASEP) [4, 23, 37] describes the flow of particles in $n$ sites along a 1D lattice. Each site may be empty or contain a particle, and particles hop randomly in a unidirectional manner, that is, from site $i$ to site $i+1$. Importantly, a particle can only hop into an empty site (simple exclusion), yielding an intricate coupling between the particles. In particular, if a particle is “stuck” at a site for an extended time, then other particles will accumulate in the sites behind it, thus generating a “traffic jam”. TASEP has found many useful applications, including modeling the flow of ribosomes along mRNA [49] and other intracellular transport processes [5], pedestrian and vehicular traffic [14], the movement of ants, and numerous other natural and artificial phenomena [37]. However, rigorous analysis of TASEP in the case where the transition rates between sites are different is non-trivial. Existing results, for example, on contraction in TASEP [10, 11] do not provide the level of detail that deterministic models can provide.
The ribosome flow model (RFM) is a deterministic model for the flow of ribosomes along the mRNA molecule. It can be derived as a dynamic mean-field approximation of TASEP [35], and is composed of $n$ non-linear first-order ODEs. The RFM can also be derived via a special finite volume spatial discretization of widely used hyperbolic PDE flow models [41]. The RFM can be interpreted as a compartmental chemical reaction network (CRN) with transition rates that depend on the amount of particles and free space in various compartments [45], and can also be represented as a port-Hamiltonian system [43].
The RFM is highly amenable to analysis using tools from systems and control theory. The RFM is a contractive system [1], a cooperative system [27], and also a totally positive differential system [26].
The RFM and its variants have been used extensively to model and analyze the flow of ribosomes along the mRNA during translation (see, e.g., [31, 32, 3, 48, 34, 47, 24, 17, 18, 29]). In particular, Ref. [28] studied the RFM with positive feedback from the production rate to the initiation rate, as a model of ribosome recycling. It was shown that the dynamic behavior of this model is simple: the closed-loop system admits a unique equilibrium that is globally asymptotically stable.
Interconnected RFM networks have been used to model and analyze large-scale translation in the cell, and in particular the effect of competition for shared resources such as ribosomes and tRNA molecules [12, 33, 30, 19, 20]. Models based on networks of RFMs have also been validated experimentally. It was shown that such a model can predict the density of ribosomes along different mRNAs, the protein levels of different genes, and can even be used to generate ribosomal traffic jams (see, for example, [50, 35, 13]).
In the case where the entry, exit, and elongation rates of the RFM are time-varying and jointly periodic, with a common period $T$, the dynamics admit a unique $T$-periodic solution that is globally exponentially stable [25]. Interpreting the rates as a periodic excitation (e.g., due to the periodic cell-cycle process) implies that the state variables in the RFM, and thus also the protein production rate, entrain to the periodic excitation (see also [36] for some related considerations).
Gene expression is tightly regulated, and negative feedback loops are common motifs that can drive the protein production rate to a desired level, as well as reduce expression noise and intercellular variability in protein levels [38]. For example, bacteria use RNA binding proteins (RBPs) as regulators of many cellular processes, including translation, and most RBPs affect translation at the level of translation initiation. Many translation-related RBPs inhibit their own mRNA [16, 2]. Note also that in the context of mRNA transcription, it is known that a negative feedback mechanism induces oscillations in the notch effector Hes1 [15].
Here, we analyze for the first time the RFM with a negative feedback interconnection from the protein production rate to the initiation rate. We show that:
  1. 1.
    The closed-loop system admits a unique equilibrium point $e$. This corresponds to a dynamic steady-state where the ribosome flow into and out of each site are equal, and thus to a constant protein production rate;
  2. 2.
    Under a simple and easy to verify sufficient condition on the local (in)stability of the system at $e$, the closed-loop system admits a periodic solution, and we provide an explicit description of a set $S$ such that any solution emanating from $S$ converges to a periodic solution.
These results suggest that negative feedback in translation can provide a simple mechanism for producing self-sustained (i.e., unforced) oscillations in protein production. It is known that translational control of cyclin B1 mRNA might play an essential role in the periodic cell-cycle process [9]. This is attributed to oscillations in the eIF4E binding factor, whose expression is cell cycle regulated. However, our results suggest that the simpler mechanism of negative feedback in translation can also lead to periodic oscillations in protein production. This agrees with the findings in [42] that used ribosome profiling to conclude that gene-specific translational repression controls the mitotic proteome, complementing post-translational mechanisms for inactivating protein function.
The fact that negative feedback may lead to oscillations is not surprising, but rigorously proving this in $n$-dimensional non-linear systems like the RFM is non-trivial. Note also that property 2) above implies that the closed-loop system is not contractive nor cooperative. Our analysis is based on showing that the closed-loop system is a strongly 2-cooperative system [46, 21].
The remainder of this paper is organized as follows. The next section describes the RFM with negative feedback. Section III describes the main analysis results. The last section concludes. The proofs of the main results are placed in the Appendix.
II The model
The dynamic equations of the ribosome flow model (RFM) with $n$ sites are:
$\dot{x}_{1}(t)$ $=\lambda_{0}(1-x_{1}(t))-\lambda_{1}x_{1}(t)(1-x_{2}(t)),$
$\dot{x}_{2}(t)$ $=\lambda_{1}x_{1}(t)(1-x_{2}(t))-\lambda_{2}(t)x_{2}(t)(1-x_{3}(t)),$
$\vdots$ (1)
$\dot{x}_{n}(t)$ $=\lambda_{n-1}x_{n-1}(t)(1-x_{n}(t))-\lambda_{n}x_{n}(t),$
where $x_{i}(t)\in[0,1]$ is the ribosome density levels at site $i$ at time $t$, and $\lambda_{i}>0,i=0,...,n$ are the translation rates from site $i$ to site $i+1$ at time $t$. To explain this non-linear model, consider the general equation
$\dot{x}_{i}(t)=\lambda_{i-1}x_{i-1}(t)(1-x_{i}(t))-\lambda_{i}x_{i}(t)(1-x_{i+1}(t)).$ (2)
This can be explained as follows. The term $\lambda_{i-1}x_{i-1}(1-x_{i})$ is the flow rate from site $i-1$ to site $i$. This depends on: (1) the parameter $\lambda_{i-1}$ that models “how easy it is” for a ribosome to move from site $i-1$ to site $i$ (e.g., due to the abundance of the cognate tRNA molecules); (2) the density $x_{i-1}$ at site $i-1$; and (3) the “free space” $1-x_{i}$ at site $i$. This is a “soft” version of the simple exclusion principle: as the density in site $i$ increases, the flow into this site decreases. Similarly, the second term on the right-hand side of (2) is the flow rate from site $i$ to site $i+1$.
To study large-scale translation in the cell using interconnections of RFMs, Ref. [28] introduced the ribosome flow model with input and output (RFMIO). This is derived by adding a scalar input function and a scalar output function to the RFM. The input $u(t)\geq 0$ controls the initiation rate to the mRNA at time $t$, and the output $y(t)$ is the protein production rate at time $t$. Thus, the RFMIO is described by
$\dot{x}_{1}(t)$ $=\lambda_{0}u(t)(1-x_{1})-\lambda_{1}x_{1}(t)(1-x_{2}(t)),$
$\dot{x}_{2}(t)$ $=\lambda_{1}x_{1}(t)(1-x_{2}(t))-\lambda_{2}x_{2}(t)(1-x_{3}(t)),$
$\vdots$ (3)
$\dot{x}_{n}(t)$ $=\lambda_{n-1}x_{n-1}(t)(1-x_{n}(t))-\lambda_{n}x_{n}(t),$
$y(t)$ $=\lambda_{n}x_{n}(t).$
RFM with negative feedback
Here, we consider for the first time the RFMIO (II) where the input $u(t)$ is given by
$u(t)=kg(y(t)),$ (4)
with $k>0$ and $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a continuously differentiable function such that $g^{\prime}(z)=\frac{dg(z)}{dz}<0$ for any $z>0$. Thus, $g$ is a monotonically decreasing function. We refer to $k$ as the feedback gain. This models negative feedback from the protein production rate $y(t)$ to the RFM initiation rate (that is controlled by the input $u(t)$).
The closed-loop system that arises from (II) by eliminating $y$ and substituting $u$ from (4) is
$\dot{x}_{1}(t)$ $=\lambda_{0}kg(\lambda_{n}x_{n}(t))(1-x_{1}(t))-\lambda_{1}x_{1}(t)(1-x_{2}(t)),$
$\dot{x}_{2}(t)$ $=\lambda_{1}x_{1}(t)(1-x_{2}(t))-\lambda_{2}x_{2}(t)(1-x_{3}(t)),$
$\vdots$ (5)
$\dot{x}_{n}(t)$ $=\lambda_{n-1}x_{n-1}(t)(1-x_{n}(t))-\lambda_{n}x_{n}(t).$
Let $f:[0,1]^{n}\to\mathbb{R}^{n}$ be the vector field on the right-hand side of (II), so that we can write the closed-loop system as
$\dot{x}=f(x),$
and let
$J(x):=\frac{\partial}{\partial x}f(x)$
be the Jacobian of the vector field $f$. For an initial condition $x(0)=a$, let $x(t,a)$ denote the solution of the closed-loop system at time $t\geq 0$.
III Main results
We can now state our main results. The proofs of these results are given in the Appendix. Our first result guarantees in particular that the unit cube $[0,1]^{n}$ is a forward invariant set for the closed-loop system. In other words, if we initialize the state-variables $x_{i}(0)$ at time zero with admissible density values, that is, all values are in the range $[0,1]$ then they maintain admissible values for all $t>0$.
Proposition 1.
For any initial condition $a\in[0,1]^{n}$, we have that $x(t,a)\in(0,1)^{n}$ for all $t>0$.
Our second main result shows that the closed-loop system admits a unique equilibrium $e\in[0,1]^{n}$. At this equilibrium, that is, when $x_{i}(t)=e_{i}$ for all $i$, the flow into each site is equal to the flow out of the site, and thus the density at each site remains unchanged. Furthermore, for any feedback gain $k>0$ sufficiently small the equilibrium is globally asymptotically stable, that is, the density along the mRNA will converge to $e$ from any initial condition.
Proposition 2.
The closed-loop system described by (II) and (4) admits a unique equilibrium $e$ in the unit cube, and $e\in(0,1)^{n}$. If the feedback gain $k>0$ is sufficiently small then any solution of the closed-loop system converges to $e$.
Our third main result provides a simple condition guaranteeing that the closed-loop system admits a (non-trivial) periodic orbit. To state it, we recall the notion of the weak number of sign variations in a vector [8].
If $z\in\mathbb{R}^{n}$ has no zero entries then the number of sign variation in $z$, denoted $\sigma(z)$, is the number of indexes $i$ such that $z_{i}z_{i+1}<0$. For example, for $z=\begin{bmatrix}1&-2&4\end{bmatrix}^{\top}$, we have $z_{1}z_{2}<0$ and $z_{2}z_{3}<0$, so $\sigma(z)=2$. More generally, for any vector $z\in\mathbb{R}^{n}$, with $z\not=0$, the weak number of sign variations in $z$ is defined as $s^{-}(z):=\sigma(\tilde{z})$, where $\tilde{z}$ is the vector obtained from $z$ after deleting all its zero entries. For example, for $z=\begin{bmatrix}0&1&-2&0&4\end{bmatrix}^{\top}$, we have $\tilde{z}=\begin{bmatrix}1&-2&&4\end{bmatrix}^{\top}$, so $s^{-}(z)=\sigma(\tilde{z})=2$. For the zero vector, define $s^{-}(0):=0$. Note that for any $z\in\mathbb{R}^{n}$ and any non-zero scalar $c$, we have $s^{-}(cz)=s^{-}(z)$.
An eigenvalue $\lambda$ of a matrix $A\in\mathbb{R}^{n\times n}$ is called unstable if the real part of $\lambda$ is positive.
Theorem 3.
Let $e\in(0,1)^{n}$ denote the unique equilibrium of the closed-loop system described by (II) and (4). If $J(e)=\frac{\partial}{\partial x}f(e)$ has at least two unstable eigenvalues then the following property holds. For any initial condition $a\in[0,1]^{n}$ such that $a\not=e$ and $s^{-}(a-e)\leq 1$ the corresponding solution $x(t,a)$ converges to a non-trivial periodic orbit.
Note that since $e\in(0,1)^{n}$, we have $s^{-}(0-e)=0$ implying that if $J(e)$ has at least two unstable eigenvalues then the solution $x(t,0)$ always converges to a non-trivial periodic orbit. The zero initial condition corresponds in our model to no ribosomes along the mRNA molecule, which is the case for a newly synthesized mRNA.
Example 1.
Consider the closed-loop system with $n=3$, $\lambda_{0}=12$, $\lambda_{1}=6$, $\lambda_{2}=5$, $\lambda_{3}=8$, and $u(t)=kg(y(t))$, where $g(s)=\frac{1}{1+s^{20}}$. When $k=0.1$, a calculation shows that $e=\begin{bmatrix}0.1855&0.2004&0.1113\end{bmatrix}^{\top}$ (all numerical values are to 4-digit accuracy), and the eigenvalues of $J(e)$ are
$-14.9362,\;-2.7564+6.0039i,\;-2.7564-6.0039i.$
Note that these all have negative real parts. Fig. 1 depicts several trajectories of the system, and it may be seen that they all converge to $e$.
When $k=1$, a calculation shows that $e=\begin{bmatrix}0.2472&0.2560&0.1380\end{bmatrix}^{\top},$ and the eigenvalues of $J(e)$ are
[…]
so $J(e)$ admits two unstable eigenvalues. Fig. 2 depicts the solution $x(t,a)$ of the closed-loop system emanating from the initial condition $a=0$. Note that $s^{-}(a-e)=0$. It may be seen that this solution does not converge to $e$, but rather converges to a periodic orbit.
Note that Theorem 3 holds for the closed-loop system with any dimension $n$. In Example 1 we considered the special case $n=3$ only because in this case it is easy to plot the solutions of the system.
Figure 1: Several trajectories of the closed-loop system in Example 1 with $k=0.1$. The initial conditions are marked with $\circ$ and the equilibrium point $e$ is marked by *.
Figure 2: A trajectory of the closed-loop system in Example 1 with $k=1$. The equilibrium point $e$ is marked by *.
IV Discussion
We considered a mathematical model for mRNA translation with a negative feedback connection from the protein production rate to the initiation rate. This may model for example the effect of RBPs that inhibit their own mRNA, as well as other gene expression regulation mechanisms like the production of proteins that prevent the assembly of the translation initiation complex. More generally, negative regulation of mRNA translation can be used to fine tune protein synthesis, maintaining cellular homeostasis and adaptivity to environmental changes or internal signals.
Our results show that the closed-loop system admits a unique steady-state $e$, but if this steady state is locally unstable (more specifically, if the Jacobian of the vector field at $e$ admits at least two unstable eigenvalues) then non-trivial oscillations appear. These correspond to a periodic pattern of protein production. This may also generate an internal clock that can regulate timing in various cellular processes.
It is interesting to note that Castelo-Szekely et al. [6] have found that tissue specificity in rhythmic gene expression extends to the translatome and contributes to define the identities, the phases and the expression levels of rhythmic protein biosynthesis. Also, it is known that translation oscillates during the periodic cell-cycle program (see, e.g., [42]), and our results suggest that this may be generated by a simple negative feedback mechanism.
More work is needed in order to analyze the properties of the oscillatory solutions in our model e.g. their amplitude and frequency, and on how these depend on the various parameters of the system.
Acknowledgments
We thank Alexander Ovseevich for helpful comments.
V Appendix: Proofs
Proof of Proposition 1. Consider the closed-loop system (II). Fix an initial condition $x(0)=a\in[0,1]^{n}$. Suppose that there exist $i\in\{1,\dots,n\}$ and $t\geq 0$ such that $x_{i}(t,a)=0$. Then there exists a minimal index $j$ such that $x_{j}(t,a)=0$. If $j=1$ then
$\dot{x}_{1}(t)=\lambda_{0}kg(\lambda_{n}x_{n}(t))>0.$
If $j\in\{2,\dots,n-1\}$ then
$\dot{x}_{j}(t)=\lambda_{j-1}x_{j-1}(t)>0.$
This implies that $x_{\ell}(t)>0$ for all $\ell\in\{1,\dots,n\}$ and all $t>0$. A similar argument shows that $x_{\ell}(t)<1$ for all $\ell\in\{1,\dots,n\}$ and all $t>0$, and this completes the proof.
Proof of Proposition 2. Since $[0,1]^{n}$ is convex and compact, it follows from Proposition 1 and Brouwer’s fixed point theorem [22] that the closed-loop system admits at least one equilibrium $e\in[0,1]^{n}$. Using Proposition 1 implies that in fact $e\in(0,1)^{n}$. By (II), the equilibrium satisfies
$\lambda_{0}kg(\lambda_{n}e_{n})(1-e_{1})$ $=\lambda_{1}e_{1}(1-e_{2})$
$=\lambda_{2}e_{2}(1-e_{3})$
$\vdots$ (6)
$=\lambda_{n-2}e_{n-2}(1-e_{n-1})$
$=\lambda_{n-1}e_{n-1}(1-e_{n})$
$=\lambda_{n}e_{n}.$
These equations imply in particular that $e_{n}$ uniquely determines $e_{i}$ for all $i$. In order to prove uniqueness of the equilibrium, seeking a contradiction, suppose that $p,q\in(0,1)^{n}$ are two different equilibrium points. Then $p_{n}\not=q_{n}$, and we may assume that $p_{n}>q_{n}$. By (V), $e_{n-1}=\frac{\lambda_{n}e_{n}}{\lambda_{n-1}(1-e_{n})}$, so $p_{n-1}>q_{n-1}$. Eq. (V) also gives $e_{n-2}=\frac{\lambda_{n}e_{n}}{\lambda_{n-2}(1-e_{n-1})}$, so $p_{n-2}>q_{n-2}$. Continuing in this way, we find that $p_{i}>q_{i}$ for all $i$. Now (V) and the fact that $g$ is decreasing give
$\lambda_{n}p_{n}$ $=\lambda_{0}kg(\lambda_{n}p_{n})(1-p_{1})$
$<\lambda_{0}kg(\lambda_{n}q_{n})(1-q_{1})$
$=\lambda_{n}q_{n},$
so $p_{n}<q_{n}$. This contradiction shows that the equilibrium is unique.
The Jacobian of the closed-loop system is
$J(x)=\left[\begin{smallmatrix}-\lambda_{0}kg(\lambda_{n}x_{n})-\lambda_{1}(1-x_{2})&\lambda_{1}x_{1}&0&0&\dots&0&\lambda_{0}\lambda_{n}(1-x_{1})kg^{\prime}(\lambda_{n}x_{n})\\ \lambda_{1}(1-x_{2})&-\lambda_{1}x_{1}-\lambda_{2}(1-x_{3})&\lambda_{2}x_{2}&0&\dots&0&0\\ &&\vdots\\ 0&0&0&0&\dots&\lambda_{n-1}(1-x_{n})&-\lambda_{n-1}x_{n-1}-\lambda_{n}\end{smallmatrix}\right].$ (7)
Let $s_{i}(x)$ denote the sum of the entries in column $i$ of $J(x)$, with the off-diagonal terms taken with absolute value. Then
$s_{1}(x)$ $=-\lambda_{0}kg(\lambda_{n}x_{n})\leq-\lambda_{0}kg(\lambda_{n}),$
$s_{2}(x)$ $=\dots=s_{n-1}(x)=0,$
$s_{n}(x)$ $=-\lambda_{n}+\lambda_{0}\lambda_{n}(1-x_{1})k|g^{\prime}(\lambda_{n}x_{n})|\leq-\lambda_{n}-\lambda_{0}\lambda_{n}k\min_{r\in[0,1]}g^{\prime}(\lambda_{n}r).$
For any $k>0$ sufficiently small, we have that $s_{1}(x),s_{n}(x)<0$ for all $x\in[0,1]^{n}$. Recall that the matrix measure induced by the $L_{1}$ norm is $\mu_{1}(J(x))=\max\{s_{1}(x),\dots,s_{n}(x)\}$ (see, e.g., [1]). Using a scaled $L_{1}$ norm, as in [25], yields that the system is contractive. This implies that $e$ is globally asymptotically stable.
Proof of Theorem 3. Since $g^{\prime}(z)<0$ for all $z\geq 0$, Eq. (7) implies that $J(x)$ has the sign pattern of a 2-cooperative system [46] for all $x\in[0,1]^{n}$, and it is irreducible for all $x\in(0,1)^{n}$. Using Theorem 2 in [21] completes the proof of Theorem 3.