### v2

# Negative feedback and oscillations in a model for mRNA translation

Abstract
The ribosome flow model (RFM) is a phenomenological model for the unidirectional flow of particles along a 1D chain of $n$ sites. The RFM has been extensively used to study the dynamics of ribosome flow along a single mRNA molecule during translation. In this case, the particles model ribosomes and each site corresponds to a consecutive group of codons. Networks of interconnected RFMs have been used to model and analyze large-scale translation in the cell and, in particular, the effects of competition for shared resources. Here, we analyze the RFM with a negative feedback connection from the protein production rate to the initiation rate. This models, for example, the production of proteins that inhibit the translation of their own mRNA. The RFM with negative feedback is a 2-cooperative dynamic system, i.e. its flow maps the set of vectors with up to one sign variation to itself. Using tools from the theory of 2-cooperative dynamical systems, we provide a simple condition guaranteeing that the closed-loop system admits at least one non-trivial periodic solution. When this condition holds, we also explicitly characterize a large set of initial conditions such that any solution emanating from this set converges to a non-trivial periodic solution. Such a solution corresponds to a periodic pattern of ribosome densities along the mRNA, and to a periodic pattern of protein production.
Index Terms:
mRNA translation, periodic gene expression, 2-cooperative systems.
I Introduction
Living cells produce proteins using RNA molecules as templates. Three consecutive nucleotides in the mRNA form a codon that encodes for a specific amino-acid. Complex molecular machines, called ribosomes, scan the mRNA sequentially, codon by codon, and generate a chain of corresponding amino-acids that eventually becomes the coded protein. Every ribosome that reaches the stop codon releases the chain of amino acids, so the rate of ribosomes leaving the chain is the protein production rate. To pipeline this process, several ribosomes may decode the same mRNA molecule simultaneously, thus generating a polysome.
The dynamics of ribosome flow along the mRNA is determined by various bio-physical properties including the mRNA sequence, abundance of decoding factors, and the interactions between ribosomes. Indeed, since every ribosome must follow the mRNA in a sequential manner, a slow-moving ribosome may lead to the generation of “traffic jams” of ribosomes behind it. Such traffic jams have been implicated with various diseases including Huntington disease [50],  [49].
A fundamental problem in biology and medicine is understanding the regulation and dynamics of ribosome flow and the protein production rate. Rigorous analysis of ribosome flow is also important in scientific fields that include natural or artificial manipulations of the translation machinery such as biotechnology, synthetic biology, and mRNA viruses.
I-A Modeling ribosome flow
Computational modeling of ribosome flow is attracting considerable interest [54, 11, 61]. This is motivated also by new experimental methods, like ribosome profiling [23], that provide considerable information on which mRNAs are being actively translated in a cell at a given time.
The totally asymmetric simple exclusion process (TASEP) [8, 31, 47] describes the flow of particles along a 1D lattice with $n$ sites. Each site may be empty or contain a particle, and particles hop randomly in a unidirectional manner, that is, from site $i$ to site $i+1$. Importantly, a particle can only hop into an empty site (simple exclusion), yielding an intricate indirect coupling between the particles. In particular, if a particle is “stuck” at a site for an extended period of time, then other particles will accumulate in the sites behind it, thus generating a “traffic jam”. TASEP has found many useful applications, including modeling the flow of ribosomes along mRNA [59] and other intracellular transport processes [9], pedestrian and vehicular traffic [19], the movement of ants, and numerous other natural and artificial phenomena [47]. However, rigorous analysis of TASEP in the case where the transition rates between sites are different is non-trivial. Existing results, for example, on contraction in TASEP [15, 16] do not provide the level of detail that deterministic models can provide.
The ribosome flow model (RFM) is a deterministic model for the flow of ribosomes along the mRNA molecule. It can be derived as a dynamic mean-field approximation of TASEP [45], and is composed of $n$ non-linear first-order ODEs. The RFM can also be derived via a special finite volume spatial discretization of widely used hyperbolic PDE flow models [51]. The RFM can be interpreted as a compartmental chemical reaction network (CRN) with transition rates that depend on the amount of particles and free space in various compartments [55], and can also be represented as a port-Hamiltonian system [53].
The RFM is highly amenable to analysis using tools from systems and control theory. The RFM is a contractive system [2], a cooperative system [36], and also a totally positive differential system [35].
The RFM and its variants have been used extensively to model and analyze the flow of ribosomes along the mRNA during translation (see, e.g., [41, 42, 5, 58, 44, 57, 33, 24, 25, 38]). In particular, Ref. [37] studied the RFM with positive feedback from the production rate to the initiation rate, as a model of ribosome recycling. It was shown that the dynamic behavior of this model is simple: the closed-loop system admits a unique equilibrium that is globally asymptotically stable.
Interconnected RFM networks have been used to model and analyze large-scale translation in the cell, and in particular the effect of competition for shared resources such as ribosomes and tRNA molecules [17, 43, 39, 26, 28]. Models based on networks of RFMs have also been validated experimentally. It was shown that such a model can predict the density of ribosomes along different mRNAs, the protein levels of different genes, and can even be used to generate ribosomal traffic jams (see, for example, [60, 45, 18]).
In the case where the entry, exit, and elongation rates in the RFM are time-varying and jointly periodic, with a common period $T$, the dynamics admit a unique $T$-periodic solution that is globally exponentially stable [34]. Interpreting the rates as a periodic excitation (e.g., due to the periodic cell-cycle process) implies that the state variables in the RFM, and thus also the protein production rate, entrain to the periodic excitation (see also [46] for some related considerations).
Gene expression is tightly regulated, and negative feedback loops are common motifs that can drive the protein production rate to a desired level, as well as reduce expression noise and intercellular variability in protein levels [48].
Gene expression is highly regulated to ensure that the right proteins are made in the right places within the cell. Here, we are interested in analyzing the effects of negative feedback in the level of translation. For example, bacteria use RNA binding proteins (RBPs) as regulators of many cellular processes, including translation, and most RBPs affect translation at the level of translation initiation. Many translation-related RBPs inhibit their own mRNA [21, 3].
To study this process, we analyze for the first time the RFM with a negative feedback interconnection from the protein production rate to the initiation rate. We show that:
  1. 1.
    The closed-loop system admits a unique equilibrium point $e$. This corresponds to a steady-state where the ribosome flow into and out of each site are equal, and thus to a constant protein production rate;
  2. 2.
    Under a simple and easy to verify sufficient condition on the local (in)stability of the system at $e$, the closed-loop system admits a periodic solution, and we provide an explicit description of a set $S$ such that any solution emanating from $S$ converges to a (non-trivial) periodic solution.
These results suggest that negative feedback in the level of translation only can provide a simple cellular mechanism for producing self-sustained (i.e., unforced) oscillations in protein production. This agrees with the findings in [52] that used ribosome profiling to conclude that gene-specific translational repression controls the mitotic proteome, complementing post-translational mechanisms for inactivating protein function.
The fact that negative feedback may lead to oscillations is not surprising, but rigorously proving this in $n$-dimensional non-linear systems like the RFM is non-trivial. Note also that property 2) above implies that the closed-loop system is not contractive nor cooperative. Our analysis is based on showing that the closed-loop system is a strongly 2-cooperative system [56, 29]. Roughly speaking, strongly 2-cooperative dynamic systems map the set of vectors with up to one sign variation to itself. Such systems have a Poincaré-Bendixson property, which implies that bounded solutions which do not converge to an equilibrium converge to a periodic orbit. Ref. [29] derived specific conditions on the Jacobian at the equilibrium such that periodic orbits exist, and in this case also described a set of initial conditions that lead to a periodic solution.
I-B Negative feedback and cellular oscillations
The emergence of sustained oscillations (via convergence to periodic orbits) in high-dimensional nonlinear dynamical systems is a highly non-trivial question with important applications in systems biology, including the understanding of bio-molecular oscillators ruling cell life-cycle and metabolism, as well as circadian rhythms in hormone secretion, body temperature and metabolic functions [14].
Many cellular oscillatory processes, including certain circadian oscillators, are based on the transcription-translation feedback loop (TTFL), that includes both positive and negative feedback loops [22]. The recurring motifs of negative feedback loops, positive feedback loops, and reciprocal regulation play a crucial role in timing and regulating cell-cycle oscillations, and it is known that sufficiently strong negative feedback induces oscillations [13]. For a very readable account of the role of feedback loops in biological oscillators, see [1, Chapter 6].
The p53–MDM2 loop is a well-known negative feedback circuit that plays a crucial role in regulating cell fate decisions. p53 activates the transcription of many genes, including mdm2. MDM2 inhibits the transcriptional activity of p53 and increases degradation of p53. This mechanism yields oscillations or pulses in p53 levels in response to DNA damage [4, 7, 6].
A negative feedback mechanism in transcription induces oscillations in the notch effector Hes1 [20]. A “minimal” mathematical model for this phenomena, that includes two ordinary differential equations and a time delay in the negative feedback loop, was derived in [27].
An important goal in synthetic biology is building genetic oscillators that can be used for correctly timing and synchronizing various processes. Elowitz and Leibler [32] designed an artificial oscillator in E. coli consisting of a loop with three transcriptional repressors. Negative feedback loops are known to be a key design principle for synthetic biological oscillators [40], but constructing reliable biomolecular oscillators with tunable amplitude and phase is still an open challenge.
The model that we explore here suggests that oscillations can also be generated by direct negative feedback from the produced protein to the mRNA initiation rate.
The remainder of this paper is organized as follows. The next section describes the RFM with negative feedback. Section III describes the main analysis results. The last section concludes. The proofs of the main results are placed in the Appendix 1.
II The model
The dynamic equations of the ribosome flow model (RFM) with $n$ sites are:
$\dot{x}_{1}(t)$ $=\lambda_{0}(1-x_{1}(t))-\lambda_{1}x_{1}(t)(1-x_{2}(t)),$
$\dot{x}_{2}(t)$ $=\lambda_{1}x_{1}(t)(1-x_{2}(t))-\lambda_{2}(t)x_{2}(t)(1-x_{3}(t)),$
$\vdots$ (1)
$\dot{x}_{n}(t)$ $=\lambda_{n-1}x_{n-1}(t)(1-x_{n}(t))-\lambda_{n}x_{n}(t),$
where $x_{i}(t)\in[0,1]$ is the ribosome density levels at site $i$ at time $t$, and $\lambda_{i}>0,i=0,...,n$ are the translation rates from site $i$ to site $i+1$ at time $t$. Note that the $x_{i}$s are dimensionless and the $\lambda_{i}$s have units of $1/\text{time}$.
To explain this non-linear model, consider the general equation
$\dot{x}_{i}(t)=\lambda_{i-1}x_{i-1}(t)(1-x_{i}(t))-\lambda_{i}x_{i}(t)(1-x_{i+1}(t)).$ (2)
This can be explained as follows. The term $\lambda_{i-1}x_{i-1}(1-x_{i})$ is the flow rate from site i1i-1 to site $i$. This depends on: (1) the parameter $\lambda_{i-1}$ that models “how easy it is” for a ribosome to move from site i1i-1 to site $i$ (e.g., due to the abundance of the cognate tRNA molecules); (2) the density $x_{i-1}$ at site i1i-1; and (3) the “free space” $1-x_{i}$ at site $i$. This is a “soft” version of the simple exclusion principle: as the density in site $i$ increases, the flow into this site decreases. Similarly, the second term on the right-hand side of (2) is the flow rate from site $i$ to site $i+1$.
To study large-scale translation in the cell using interconnections of RFMs, Ref. [37] introduced the ribosome flow model with input and output (RFMIO). This is derived by adding a scalar input function and a scalar output function to the RFM. The input $u(t)\geq 0$ controls the initiation rate to the mRNA at time $t$, and the output $y(t)$ is the protein production rate at time $t$. Thus, the RFMIO is described by
$\dot{x}_{1}(t)$ $=\lambda_{0}(1-x_{1}(t))u(t)-\lambda_{1}x_{1}(t)(1-x_{2}(t)),$
$\dot{x}_{2}(t)$ $=\lambda_{1}x_{1}(t)(1-x_{2}(t))-\lambda_{2}x_{2}(t)(1-x_{3}(t)),$
$\vdots$ (3)
$\dot{x}_{n}(t)$ $=\lambda_{n-1}x_{n-1}(t)(1-x_{n}(t))-\lambda_{n}x_{n}(t),$
$y(t)$ $=\lambda_{n}x_{n}(t).$
RFM with negative feedback
Here, we consider for the first time the RFMIO (II) where the input $u(t)$ is given by
$u(t)=kg(y(t)),$ (4)
with $k>0$ and $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a continuously differentiable function such that $g^{\prime}(z)=\frac{dg(z)}{dz}<0$ for any $z>0$. Thus, $g$ is a monotonically decreasing function. We refer to $k$ as the feedback gain. This models negative feedback directly from the protein production rate $y(t)$ to the RFM initiation rate that is controlled by the input $u(t)$ (see Fig. 1).
Note that (4) captures many negative feedback functions used in models from systems biology and in particular gene regulation. For example, g(y)=ayg(y)=-ay, with $a>0$ (negative linear feedback), $g(y)=e^{-ay}$, with $a>0$ (exponential decay), $g(y)=\frac{1}{a+y}$, with $a>0$ (Michaelis-Menten inhibition), $g(y)=\frac{1}{1+(y/a)^{\ell}}$, with $a,\ell>0$ (Hill type inhibition), $g(y)=1-\tanh(ay)$, with $a>0$, and more.
[…]
Figure 1: RFMIO with feedback $u=kg(y)$.
The closed-loop system that arises from (II) by eliminating $y$ and substituting $u$ from (4) is
$\dot{x}_{1}(t)$ $=\lambda_{0}kg(\lambda_{n}x_{n}(t))(1-x_{1}(t))-\lambda_{1}x_{1}(t)(1-x_{2}(t)),$
$\dot{x}_{2}(t)$ $=\lambda_{1}x_{1}(t)(1-x_{2}(t))-\lambda_{2}x_{2}(t)(1-x_{3}(t)),$
$\vdots$ (5)
$\dot{x}_{n}(t)$ $=\lambda_{n-1}x_{n-1}(t)(1-x_{n}(t))-\lambda_{n}x_{n}(t).$
Let $f:[0,1]^{n}\to\mathbb{R}^{n}$ be the vector field on the right-hand side of (II), so that we can write the closed-loop system as
$\dot{x}=f(x),$
and let
$J(x):=\frac{\partial}{\partial x}f(x)$
be the Jacobian of the vector field $f$. For an initial condition $x(0)=a$, let $x(t,a)$ denote the solution of the closed-loop system at time $t\geq 0$.
III Main results
We can now state our main results. The proofs of these results are given in Appendix 1. Our first result guarantees in particular that the unit cube $0,1]^{n}$ is a forward invariant set for the closed-loop system. In other words, if we initialize the state-variables $x_{i}(0)$ at time zero with admissible density values, that is, all values are in the range $[0,1]$ then they maintain admissible values for all $t>0$.
Proposition 1.
For any initial condition $a\in[0,1]^{n}$, we have that $x(t,a)\in(0,1)^{n}$ for all $t>0$.
Our second main result shows that the closed-loop system admits a unique equilibrium $e\in[0,1]^{n}$. At this equilibrium, that is, when $x_{i}(t)=e_{i}$ for all $i$, the flow into each site is equal to the flow out of the site, and thus the density at each site remains unchanged. Furthermore, for any feedback gain $k>0$ sufficiently small the equilibrium is globally asymptotically stable, that is, the density along the mRNA will converge to $e$ from any initial condition.
Proposition 2.
The closed-loop system described by (II) and (4) admits a unique equilibrium $e$ in the unit cube, and $e\in(0,1)^{n}$. If the feedback gain $k>0$ is sufficiently small then any solution of the closed-loop system converges to $e$.
Our third main result provides a simple condition guaranteeing that the closed-loop system admits a (non-trivial) periodic orbit. To state it, we recall the notion of the weak number of sign variations in a vector [12].
[…]
An eigenvalue $\lambda$ of a matrix $A\in\mathbb{R}^{n\times n}$ is called unstable if the real part of $\lambda$ is positive.
Theorem 3.
Let $e\in(0,1)^{n}$ denote the unique equilibrium of the closed-loop system described by (II) and (4). If $J(e)=\frac{\partial}{\partial x}f(e)$ has at least two unstable eigenvalues then the following property holds. For any initial condition $x(0)\in[0,1]^{n}$ such that $x(0)\not=e$ and $s^{-}(x(0)-e)\leq 1$ the corresponding solution $x(t,x(0))$ converges to a non-trivial periodic orbit.
Note that since $e\in(0,1)^{n}$, we have $s^{-}(0-e)=0$ implying that if $J(e)$ has at least two unstable eigenvalues then the solution $x(t,0)$ always converges to a non-trivial periodic orbit. The zero initial condition corresponds in our model to zero ribosomes along the mRNA molecule at the initial time, which is the case for a newly synthesized mRNA.
The next example demonstrates Theorem 3 using a numerical simulation. The Matlab code for generating this simulation is provided in Appendix 2.
Example 1.
Consider the closed-loop system with $n=3$, $\lambda_{0}=12$, $\lambda_{1}=6$, $\lambda_{2}=5$, $\lambda_{3}=8$, and $u(t)=kg(y(t))$, where $g(s)=\frac{1}{1+s^{20}}$. When $k=0.1$, a calculation shows that $e=\begin{bmatrix}0.1855&0.2004&0.1113\end{bmatrix}^{\top}$ (all numerical values in this paper are to 4-digit accuracy), and the eigenvalues of $J(e)$ are
14.9362,2.7564+6.0039i,2.75646.0039i.-14.9362,\;-2.7564+6.0039i,\;-2.7564-6.0039i.
Note that these all have negative real parts. Fig. 2 depicts several trajectories of the system, and it may be seen that they all converge to $e$.
When $k=1$, a calculation shows that $e=\begin{bmatrix}0.2472&0.2560&0.1380\end{bmatrix}^{\top}$, and the eigenvalues of $J(e)$ are
21.0188, 0.0075+11.9317i, 0.007511.9317i,-21.0188,\;0.0075+11.9317i,\;0.0075-11.9317i,
so $J(e)$ admits two unstable eigenvalues. Fig. 3 depicts the solution $x(t,a)$ of the closed-loop system emanating from the initial condition $a=0$. Note that $s^{-}(a-e)=0$. It may be seen that this solution does not converge to $e$, but rather converges to a periodic orbit.
Figure 2: Convergence of several trajectories of the closed-loop system in Example 1 to the equilibrium point with $k=0.1$. The initial conditions are marked with $\circ$ and the equilibrium point $e$ is marked by *.
Figure 3: Convergence of a trajectory of the closed-loop system in Example 1 to a periodic orbit with $k=1$.
The equilibrium point $e$ is marked by *.
Note that Theorem 3 holds for the closed-loop system with any dimension $n$. In Example 1 we considered the special case $n=3$ only because in this case it is easy to plot the solutions of the system.
Our theoretical results can be easily applied to RFMs with transition rates that are based on biological data. The analysis flow in such a case is demonstrated in the following example. The Matlab code for the following calculations is provided in Appendix 3.
Example 2.
Consider the RFM derived in [57] for the S. cerevisiae gene YBL025W that encodes the protein RRN10, which is related to regulation of RNA polymerase I. This gene has 145 codons (excluding the stop codon), and the corresponding RFM is based on dividing the mRNA into 6 consecutive pieces that include 24, 25, 25, 25, 25, and 21 non-overlapping codons, respectively. The first piece corresponds to the initiation stage. The corresponding RFM thus includes $n=5$ sites and six rates $\lambda_{0},\dots,\lambda_{5}$. These rates were estimated using ribo-seq data (see [57]) yielding
$\begin{bmatrix}\lambda_{0}&\dots&\lambda_{5}\end{bmatrix}=\begin{bmatrix}0.1678&0.2572&0.2758&0.2514&0.2612&0.3002\end{bmatrix}$, (6)
with units of $\text{sec}^{-1}$. Assume that we close the loop around the corresponding RFMIO with the negative feedback function
$g(y)=1-\tanh(ay),$
with $a>0$. This function takes values in $[0,1]$, and for large values of $a$ it decays quickly with $y$, resembling an on-off switch. Note that $g^{\prime}(y)=-\frac{a}{\cosh^{2}(ay)}$.
The equilibrium point $e$ satisfies the equations
$\lambda_{0}k(1-\tanh(a\lambda_{5}e_{5}))(1-e_{1})$ $=\lambda_{1}e_{1}(1-e_{2})$
$=\lambda_{2}e_{2}(1-e_{3})$
$=\lambda_{3}e_{3}(1-e_{4})$
$=\lambda_{4}e_{4}(1-e_{5})$
$=\lambda_{5}e_{5},$
with the $\lambda_{i}$s in (6). This set of equations can be solved numerically for any $k,a$, and Proposition 2 guarantees the existence of a unique solution $e\in(0,1)^{5}$. For example, for $k=10$ and $a=100$ the unique solution is
$e=\begin{bmatrix}0.1038&0.0978&0.1065&0.1003&0.0803\end{bmatrix}^{\top}$. (7)
The Jacobian of the closed-loop system at $e$ is
$J(e)=\left[\begin{smallmatrix}-\lambda_{0}kg(\lambda_{5}e_{5})-\lambda_{1}(1-e_{2})&\lambda_{1}e_{1}&0&0&\lambda_{0}\lambda_{5}(1-e_{1})kg^{\prime}(\lambda_{5}e_{5})\\ \lambda_{1}(1-e_{2})&-\lambda_{1}e_{1}-\lambda_{2}(1-e_{3})&\lambda_{2}e_{2}&0&0\\ 0&\lambda_{2}(1-e_{3})&-\lambda_{2}e_{2}-\lambda_{3}(1-e_{4})&\lambda_{3}e_{3}&0\\ 0&0&\lambda_{3}(1-e_{4})&-\lambda_{3}e_{3}-\lambda_{4}(1-e_{5})&\lambda_{4}e_{4}\\ 0&0&0&\lambda_{4}(1-e_{5})&-\lambda_{4}e_{4}-\lambda_{5}\end{smallmatrix}\right].$
Substituting the numerical values above and calculating the eigenvalues of $J(e)$ gives
0.0112+0.1894i, 0.01120.1894i,0.3855+0.3072i,0.38550.3072i,0.6301.0.0112+0.1894i,\;0.0112-0.1894i,\;-0.3855+0.3072i,\;-0.3855-0.3072i,\;-0.6301.
Thus, two eigenvalues have a positive real part and our theoretical results imply that for any initial condition $x(0)\in[0,1]^{5}$ such that $x(0)\not=e$ and $s^{-}(x(0)-e)\leq 1$ the corresponding solution $x(t,x(0))$ converges to a non-trivial periodic orbit. Fig. 4 depicts the solution $x(t,x(0))$ of the closed-loop system emanating from the initial condition $x(0)=0$. Note that $s^{-}(x(0)-e)=0$. It may be seen that this solution does not converge to $e$, but rather converges to a periodic orbit. Note that after convergence to the periodic pattern, it is straightforward to obtain from the numerical simulation features such as the amplitude and frequency of each density and of the protein production rate.
Figure 4: The state-variables $x_{1}(t)$ (dotted), $x_{2}(t)$ (dashed), and $x_{3}(t)$ (solid) as a function of $t$ in Example 2. Here $t$ is in sec. To avoid cluttering, $x_{4}(t)$ and $x_{5}(t)$ are not plotted.
IV Discussion
We considered a mathematical model for mRNA translation with a negative feedback connection from the protein production rate to the initiation rate. This may model for example the effect of RBPs that inhibit their own mRNA, as well as other gene expression regulation mechanisms like the production of proteins that prevent the assembly of the translation initiation complex. More generally, negative regulation of mRNA translation can be used to fine tune protein synthesis, maintaining cellular homeostasis and adaptivity to environmental changes or internal signals.
Our results show that the closed-loop system admits a unique steady-state $e$, but if this steady state is locally unstable (more specifically, if the Jacobian of the vector field at $e$ admits at least two unstable eigenvalues) then non-trivial oscillations appear. These correspond to a periodic pattern of protein production. This may also generate an internal clock that can regulate timing in various cellular processes.
It is interesting to note that Castelo-Szekely et al. [10] have found that tissue specificity in rhythmic gene expression extends to the translatome and contributes to define the identities, the phases and the expression levels of rhythmic protein biosynthesis. Also, it is known that translation oscillates during the periodic cell-cycle program (see, e.g., [52]), and our results suggest that this may be generated by a simple negative feedback mechanism.
More work is needed in order to analyze the properties of the oscillatory solutions in our model e.g. their amplitude and frequency, and on how these features depend on the various parameters of the system.
Our analysis is based on the fact that the RFM with negative feedback is a $2$-cooperative system. It is natural to ask if this structure is preserved when we add more bio-physical features into the RFM. We describe two possible generalizations that preserve the $2$-cooperative structure. First, adding a decay term to each density equation in the RFM, that is, changing (2) to
$\dot{x}_{i}(t)=-c_{i}x_{i}(t)+\lambda_{i-1}x_{i-1}(t)(1-x_{i}(t))-\lambda_{i}x_{i}(t)(1-x_{i+1}(t)),$
with $c_{i}\in\mathbb{R}$. Indeed, such a modification changes the diagonal terms of the Jacobian of the system, but the diagonal terms play no role in determining $2$-cooperativity (see Eq. (10) below). Of course, the $c_{i}$s may change the location of the equilibrium $e$, as well as the spectral properties of $J(e)$.
Another and perhaps more interesting possibility that preserves $2$-cooperativity is the feedback connection of two RFMIOs, as depicted in Fig. 5, with $g_{1}$ a monotonically decreasing function, and $g_{2}$ a monotonically increasing function. Such a closed-loop network may be used as a model for the interaction between two mRNAs, via the proteins that they produce, and our results allow to determine when oscillations appear in such a network.
[…]
Figure 5: Feedback interconnection of two RFMIOs. One RFMIO has dimension $n$ and rates $\lambda_{i}$, and the second has dimension $m$ and rates $\eta_{i}$. They are interconnected via two feedback functions $g_{1}$ and $g_{2}$.
Acknowledgments
We thank Alexander Ovseevich and Tamir Tuller for helpful comments. We are grateful to the anonymous reviewers and the editor for many helpful comments that helped us to improve this paper.
Appendix 1: Proofs
Proof of Proposition 1. Consider the closed-loop system (II). Fix an initial condition $x(0)=a\in[0,1]^{n}$. Suppose that there exist $i\in\{1,\dots,n\}$ and $t\geq 0$ such that $x_{i}(t,a)=0$. Then there exists a minimal index $j$ such that $x_{j}(t,a)=0$. If $j=1$ then
$\dot{x}_{1}(t)=\lambda_{0}kg(\lambda_{n}x_{n}(t))>0.$
If $j\in\{2,\dots,n-1\}$ then
$\dot{x}_{j}(t)=\lambda_{j-1}x_{j-1}(t)>0.$
This implies that $x_{\ell}(t)>0$ for all $\ell\in\{1,\dots,n\}$ and all $t>0$. A similar argument shows that $x_{\ell}(t)<1$ for all $\ell\in\{1,\dots,n\}$ and all $t>0$, and this completes the proof.
Proof of Proposition 2. Since $0,1]^{n}$ is convex and compact, it follows from Proposition 1 and Brouwer’s fixed point theorem [30] that the closed-loop system admits at least one equilibrium $e\in[0,1]^{n}$. Using Proposition 1 implies that in fact $e\in(0,1)^{n}$. By (II), the equilibrium satisfies
$\lambda_{0}kg(\lambda_{n}e_{n})(1-e_{1})$ $=\lambda_{1}e_{1}(1-e_{2})$
$=\lambda_{2}e_{2}(1-e_{3})$
$\vdots$ (8)
$=\lambda_{n-2}e_{n-2}(1-e_{n-1})$
$=\lambda_{n-1}e_{n-1}(1-e_{n})$
$=\lambda_{n}e_{n}.$
These equations imply in particular that $e_{n}$ uniquely determines $e_{i}$ for all $i$. In order to prove uniqueness of the equilibrium, seeking a contradiction, suppose that $p,q\in(0,1)^{n}$ are two different equilibrium points. Then $p_{n}\not=q_{n}$, and we may assume that $p_{n}>q_{n}$. By (Appendix 1: Proofs), $e_{n-1}=\frac{\lambda_{n}e_{n}}{\lambda_{n-1}(1-e_{n})}$, so $p_{n-1}>q_{n-1}$. Eq. (Appendix 1: Proofs) also gives $e_{n-2}=\frac{\lambda_{n}e_{n}}{\lambda_{n-2}(1-e_{n-1})}$, so $p_{n-2}>q_{n-2}$. Continuing in this way, we find that $p_{i}>q_{i}$ for all $i$. Now (Appendix 1: Proofs) and the fact that $g$ is decreasing give
$\lambda_{n}p_{n}$ $=\lambda_{0}kg(\lambda_{n}p_{n})(1-p_{1})$
$<\lambda_{0}kg(\lambda_{n}q_{n})(1-q_{1})$
$=\lambda_{n}q_{n},$
so $p_{n}<q_{n}$. This contradiction shows that the equilibrium is unique.
The Jacobian of the closed-loop system is
$J(x)=\left[\begin{smallmatrix}-\lambda_{0}kg(\lambda_{n}x_{n})-\lambda_{1}(1-x_{2})&\lambda_{1}x_{1}&0&0&\dots&0&\lambda_{0}\lambda_{n}(1-x_{1})kg^{\prime}(\lambda_{n}x_{n})\\ \lambda_{1}(1-x_{2})&-\lambda_{1}x_{1}-\lambda_{2}(1-x_{3})&\lambda_{2}x_{2}&0&\dots&0&0\\ &&\vdots\\ 0&0&0&0&\dots&\lambda_{n-1}(1-x_{n})&-\lambda_{n-1}x_{n-1}-\lambda_{n}\end{smallmatrix}\right].$ (9)
Recall that $g^{\prime}(z)<0$ for any $z>0$, so entry $(1,n)$ in $J(x)$ is non-positive. Let $s_{i}(x)$ denote the sum of the entries in column $i$ of $J(x)$, with the off-diagonal terms taken with absolute value. Then
$s_{1}(x)$ $=-\lambda_{0}kg(\lambda_{n}x_{n})\leq-\lambda_{0}kg(\lambda_{n}),$
$s_{2}(x)$ $=\dots=s_{n-1}(x)=0,$
$s_{n}(x)$ $=-\lambda_{n}+\lambda_{0}\lambda_{n}(1-x_{1})k|g^{\prime}(\lambda_{n}x_{n})|\leq-\lambda_{n}-\lambda_{0}\lambda_{n}k\min_{r\in[0,1]}g^{\prime}(\lambda_{n}r).$
For any $k>0$ sufficiently small, we have that $s_{1}(x),s_{n}(x)<0$ for all $x\in[0,1]^{n}$. Recall that the matrix measure induced by the $L_{1}$ norm is $\mu_{1}(J(x))=\max\{s_{1}(x),\dots,s_{n}(x)\}$ (see, e.g., [2]). Using a scaled $L_{1}$ norm, as in [34], yields that the system is contractive. This implies that $e$ is globally asymptotically stable.
Proof of Theorem 3. A 2-cooperative dynamic system has a Jacobian with the following sign pattern
$\bar{A}_{2}:=\left[\begin{smallmatrix}*&\geq 0&0&\dots&0&\leq 0\\ \geq 0&*&\geq 0&\dots&0&0\\ 0&\geq 0&*&\dots&0&0\\ &&\vdots\\ 0&0&0&\dots&*&\geq 0\\ \leq 0&0&0&\dots&\geq 0&*\\ \end{smallmatrix}\right]$ (10)
where ”*” means that the sign of the entry is not important. Since $g^{\prime}(z)<0$ for all $z\geq 0$, Eq. (9) implies that $J(x)$ has the sign pattern of a 2-cooperative system [56] for all $x\in[0,1]^{n}$, and it is irreducible for all $x\in(0,1)^{n}$. Using Theorem 2 in [29] completes the proof of Theorem 3.
Appendix 2: Matlab code for generating Example 1
%%%%% file negtive_RFM
global k  % controller gain
global lam0 lam1 lam2 lam3 % transition rates along the chain
lam0=12   ;lam1 =6 ; lam2=5 ; lam3= 8  ;
set(gcf,’renderer’,’painters’)
figure (1); hold off;
k=0.1; % low gain
init=[0 0 0]’;plot3(0,0,0,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
init=[1/2 1/2 1/2]’; plot3(1/2,1/2,1/2,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
init=[.1 .4 .4]’;plot3(.1,.4,.4,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
init=[.4 .2 .2]’;plot3(.4,.2,.2,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
box on;
grid on;
plot3(0.1855, 0.2004, 0.1113,’*k’); % plot the eq
 xlabel(’\$x_1\$’,Interpreter=’latex’,FontSize=18);
 ylabel(’\$x_2\$’,Interpreter=’latex’,FontSize=18);
 zlabel(’\$x_3\$’,Interpreter=’latex’,FontSize=18);
set(gcf,’renderer’,’painters’)
figure(2); hold off;
k=1;   %high gain
init=[0 0 0]’;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);hold on;
plot3(0.2472 ,0.2560 ,0.1380 ,’*k’); % plot the eq
box on
grid on
xlabel(’\$x_1\$’,Interpreter=’latex’,FontSize=18);
ylabel(’\$x_2\$’,Interpreter=’latex’,FontSize=18);
zlabel(’\$x_3\$’,Interpreter=’latex’,FontSize=18);
function ret=dyn_negative_RFM(t,x)
global k   % controller gain
global lam0 lam1 lam2 lam3 % transition rates along the chain
x1=x(1);x2=x(2);x3=x(3);
u= k*1/(1+(lam3*x3)^20 ); % the negative feedback
x1d = lam0*u*(1-x1)-lam1*x1*(1-x2);
x2d = lam1*x1*(1-x2)-lam2*x2*(1-x3);
x3d = lam2*x2*(1-x3)-lam3*x3;
ret=[x1d;x2d;x3d];
Appendix 3: Matlab code for calculating the Jacobian in Example 2
lam0=0.1678;
lam1=0.2572;
lam2=0.2758;
lam3=0.2514;
lam4=0.2612;
lam5=0.3002;
k=10;
a=100;
%% The eq equations
fun = @(x) [lam0*k* (1-tanh(a*lam5*x(5)))*(1-x(1))-lam1*x(1)*(1-x(2));
lam1*x(1)*(1-x(2))-lam2*x(2)*(1-x(3));
lam2*x(2)*(1-x(3))-lam3*x(3)*(1-x(4));
lam3*x(3)*(1-x(4))-lam4*x(4)*(1-x(5));
lam4*x(4)*(1-x(5))-lam5*x(5) ];
x0 = [1/2; 1/2; 1/2; 1/2; 1/2];% an initial guess
[x, fval, exitflag, output] = fsolve(fun, x0,options);
e=x; %%%   eq point
% calculate the Jacobain J(e)
y=lam5*e(5);
g=1-tanh(a*y) ;
gdot= -a/ (cosh(a*y))^2 ;
J=[-lam0*k*g-lam1*(1-e(2))  lam1*e(1)  0  0  lam0*lam5*(1-e(1))*k*gdot ;
   lam1*(1-e(2))  -lam1*e(1)-lam2*(1-e(3)) lam2*e(2)   0             0;
    0           lam2*(1-e(3))   -lam2*e(2)-lam3*(1-e(4)) lam3*e(3)   0;
    0     0      lam3*(1-e(4))     -lam3*e(3)-lam4*(1-e(5)) lam4*e(4) ;
    0     0             0           lam4*(1-e(5)) -lam4*e(4)-lam5    ];
eig(J)